OpenStudy (anonymous):

Is (-2, -5) a solution of the system, 3x + 2y = 4 and 2x - 5y = -29? A. yes B. no

hero (hero):

Substitute the x coordinate and the y coordinate into each equation and see if both are true. If so, then (-2,-5) should be the solution.

hero (hero):

I just tested the first equation, I get 3x + 2y = 4 3(-2) + 2(-5) = 4 -6 + -10 = 4 -16 = 4 False If one test fails, then the whole thing fails and the point is not a solution.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the answer wld b x=-2 y=5